OpenStudy (anonymous):

derivative of F = 100,000 (1+sin((2pi(t-60))/365)) @OpenStudy Feedbac…

OpenStudy (stapeleliz):

If I understand correctly, you have the following function: \[F(x)\, =\, 100,000\left( 1+\sin \left( \frac{2\pi\left(t\, -\, 60\right)}{365} \right) \right)\] To find the derivative, you need to apply the Chain Rule, working from the outside in. So copy down the 100,000, and then multiply by the derivative of what is inside, which is the "1 + sin(something)". This derivative will be "0 + cos(something)". Then multiply that by the derivative of the "(something)", which will be the (2pi)/365 times the derivative of the "t - 60".